Accelerator Update for March 5, 2008 - March 7, 2008
Wednesday March 5
The day shift began with the Tevatron (TeV) conducting studies, with no Recycler stash, but with Recycler experts conducting studies, with the Antiproton Source in an AP0 access for lithium lens work, with the Main Injector (MI), Linac, and Booster operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) waiting for the next store, with NuMI off for horn repair, and with the SciBooNE / MiniBooNE (BNB) experiments taking beam.
At 9:16 AM, the BNB beam permit tripped off. It reset.
Operations began preparing to send beam to Meson at 10:15 AM.
At 11:36 AM, the TeV suffered a sector F2 quench. At the time, Operations was setting up for a continuous injection of protons into the TeV when a kicker was accidentally fired and caused a beam on beam interaction, which caused the quench.
Operations put the TeV through a dry squeeze at 12:39 PM.
TeV experts completed an injection study and then began a stored beam study at 3:09 PM.
At 3:14 PM, Operations reported a failed power supply inside the Pelletron had disrupted readbacks. The Pelletron continued to operate.
At 3:28 PM, KRF6 tripped off. It reset.
At 4:11 PM, LRF5 tripped off. A Linac expert investigated. It reset.
At 4:24 PM, the 400MeV area tripped off for the first of three times in six minutes due to bad beam positions. It reset.
At 4:41 PM, BNB tripped off. It reset.
Operations put Pbar into standby at 4:48 PM.
Operations resumed sending beam to BNB at 5:31 PM.
Recycler studies ended at 5:55 PM.
Pbar began stacking at a reduced rate at 6:22 PM, but then stopped seven minutes later. An expert reported that he thought there might be an air bubble in the lens cooling system. He accessed the area and vented the system. Stacking resumed at 7:15 PM.
At 8:31 PM, KRF4 tripped off for the first of two times in 38 minutes. After the second trip, Operations reported that the station was having phase regulation problems. The station came back into phase, but its LCW feedback loop began to oscillate. After 30 minutes of this, operators turned the loop off and back on, which fixed the problem.
At 9:10 PM, LRF2 tripped off for the first of three times in three minutes. Operators investigated.
Operations brought LRF2 online at 10:06 PM. Operations resumed sending beam to the TeV, Pbar, and BNB.
Thursday March 6
The midnight shift began with the TeV conducting studies, with Operations monitoring the Recycler stash (29.9E10), with Pbar stacking (20.8E10), with NuMI off for horn repair, and with BNB taking beam.
At 2:38 AM, Operations reported having trouble injecting protons into the TeV. Operators contacted an expert who investigated and found a data problem on a TeV “State” front-end. This problem took most of the shift to find and resolve.
At 2:43 AM, KRF3 tripped off. It reset.
At 3:31 AM, the Linac buncher tripped off. It reset.
At 4:08 AM, an air compressor at TeV sector E2 tripped off. There was no beam in the TeV. Operations reset the breaker.
At 4:15 AM, KRF4 tripped off. It reset.
At 5:23 AM, a Booster vacuum pump (B:IP132) tripped off. It reset.
Operations put the TeV through a dry squeeze at 5:29 AM.
Operations put the TeV through another dry squeeze at 7:44 AM.
Operations reported at 8 AM that the Recycler stashed three times during the midnight shift.
Operations put the TeV through a wet squeeze at 8:02 AM.
TeV experts began an orbit smooth study at 9:04 AM.
Operations put the TeV through another wet squeeze at 11:09 AM.
TeV experts reported at 11:40 AM that they had completed all of their studies.
Operations put the TeV into shot setup at 11:52 AM.
At 12:04 PM, KRF4 tripped off for the first of three times in about 40 minutes. Operators discovered that the trips were due to a flow sensor. It reset.
At 12:07 PM, MI reported a broken vertical flying wire.
The Neutron Therapy Facility (NTF) reported at 12:31 PM that they had completed treatment on their last patient for the day.
A MI expert began work on a MI-62 kicker (KPS6N) at 12:57 PM. He finished his work about 90 minutes later.
At 1:10 PM, a MI kicker (KPS5S) had tripped off, which halted the proton injections into the TeV. An expert investigated and had it back online 43 minutes later.
At 2:09 PM, KRF3 tripped off. It reset.
At 2:17 PM, Operations halted Linac beam so that technicians could replace the coil for KRF4. Booster experts began replacing the 400 MeV area chopper tank.
Operations established store 5965 at 2:49 PM with an initial luminosity of 154.89E30.
Linac techs reported at 2:54 PM that KRF4 was back online.
Booster techs reported at 3:23 PM that the 400 MeV chopper was back online.
At 3:24 PM, a TeV corrector (T:HB44) power supply suffered a fuse fault. Operators will replace the fuse in between stores.
At 3:26 PM, Operations halted the Pbar beam because they could not see the lithium lens readbacks. An expert investigated and restored the lens readback about 30 minutes later.
At 3:30 PM, a Booster vacuum pump (IPRF23) tripped off and took down BRF13 & 14 with it. Everything reset.
At 3:48 PM, Operations halted the Pbar beam because the antiproton target had stopped rotating. Experts investigated and reported that the target was rotating, but the readbacks were not working yet.
Operations reported at 3:54 PM that operators had search and secured the following NuMI areas: MI-65, the Decay pipe, and the Absorber area.
At 4:42 PM, LRF3 tripped off. It reset.
At 5:48 PM, a Booster vacuum pump (IPRF17) tripped off and took down BRF7 & 8. Everything reset.
At 7:43 PM, Operations reported that the H- Source was sparking.
At 11:05 PM, a Booster vacuum pump (IP61) tripped off. It reset.
Operations reported at midnight that the Recycler had stashed four times during the evening shift.
Friday March 7
The midnight shift began with Operations monitoring store 5965 (50.96E30) and the stash (166.6E10), with Pbar stacking (34.4E10), with NuMI off for horn repair, and with BNB taking beam.
At 2:02 AM, MIRF2 tripped off. It reset.
At 3:22 AM, BRF10 tripped off. It reset.
At 4:27 AM, BRF11 tripped off. It reset.
Operations began preparing the NuMI area at 5:58 AM for safety system testing. The test was completed within an hour.
At 7:12 AM, Operations received an alarm for the Kautz road substation T85 transformer that indicated a low nitrogen/oil level. Operations contacted the Duty Electrician.
At 7:47 AM, BRF10 tripped off. It reset.
At 7:58 AM, KRF5 tripped off. It reset.
Operations reported at 8 AM that the Recycler had stashed four times during the midnight shift.
The Plans for Friday and the Weekend
The plans for today and the weekend are to stack, stash, and store. The Run Coordinator reported that he will allow a Preaccelerator access and a MI study today. NuMI will resume taking beam on Saturday afternoon and run through to Monday morning.
Comments and Suggestions
|last modified 03/08/2008 email Fermilab|
|Security, Privacy, Legal|